### v1

# Unfolding the Literature: A Review of Robotic Cloth Manipulation

Abstract
The realm of textiles spans clothing, households, healthcare, sports, and industrial applications. The deformable nature of these objects poses unique challenges that prior work on rigid objects cannot fully address. The increasing interest within the community in textile perception and manipulation has led to new methods that aim to address challenges in modeling, perception, and control, resulting in significant progress. However, this progress is often tailored to one specific textile or a subcategory of these textiles. To understand what restricts these methods and hinders current approaches from generalizing to a broader range of real-world textiles, this review provides an overview of the field, focusing specifically on how and to what extent textile variations are addressed in modeling, perception, benchmarking, and manipulation of textiles. We finally conclude by identifying key open problems and outlining grand challenges that will drive future advancements in the field.
doi:
10.1146/((please add article doi))
keywords:
Textiles, Deformable Object Manipulation, Generalization, Physical Properties Variations, Tasks Variations
journal: Xxxx. Xxx. Xxx. Xxx.
1 INTRODUCTION
Textile deformable objects such as clothing items or household objects like bed sheets and blankets are ubiquitous in our daily lives. Their usage spans applications from healthcare and domestic environments to the textile industry. Efforts to automate the manipulation and processing of these objects promise to enhance recycling and textile reuse while providing greater assistance to aging populations. Despite recent advances in manipulation tasks such as assistive dressing [1, 2], folding or bagging [3, 4], textile manipulation remains challenging as the deformable nature of these objects breaks fundamental assumptions in robotics such as rigidity, known dynamics models, and low dimensional state space [5]. Specifically, when forces are applied to a deformable body, they not only move the object but also change its shape. From a physics point of view, understanding how the shape changes requires knowledge about the object’s physical properties, such as stiffness or elasticity. From a perceptual point of view, properties such as shape, color, and material provide distinct signals to visual and tactile sensors. Endowing robots with skills to perceive, manipulate, and address the diversity of these textile properties presents a compelling avenue toward autonomous agents. However, current methods for textile manipulation tend to be tailored to specific objects that mirror the properties of the simulators used in their design [4, 6, 7, 8, 9].
This review consolidates recent methods and applications of deformable object manipulation, specifically textiles, highlighting the challenges associated with variations in their physical properties. We evaluate the progress made and identify key areas requiring further research to enhance the generalization and adaptive capabilities of robots in handling real-world textiles. Recent reviews have explored specific methods and applications related to this object category. Particular emphasis has been given to grasping [10] and caregiving scenarios  [11, 12]. In contrast, our review broadens the scope by analyzing generalization across different textile variations and applications. Building on the foundational work on perception, modeling, and manipulation of deformable objects [13, 14], our work offers a distinct perspective on enhancing the generalization and adaptability of perceptual and manipulation skills. Similarly to [5], we seek to identify ongoing challenges in modeling, perception, and control, further addressing the underexplored area of benchmarking.
The review is organized as follows. In Section 2, we provide the fundamentals about textiles, covering the variations of properties and tasks that will be discussed throughout the document. Section 3 reviews analytical and learning methods to model textile dynamics, highlighting their connection to textile physical properties. Section 4 identifies current approaches to perceive textile properties, whereas Section 5 covers current approaches for textile manipulations with a focus on techniques that enable generalization and adaptability to variations of properties and to what extent do current methods account for variations in physical and mechanical properties and tasks. In Section 6, we discuss currently available resources such as benchmarks and datasets, that enables evaluating this generalization, including benefits and limitations. We provide a thorough overview of application areas in Section 7. We close with a discussion about the interplay between modeling, perception, and manipulation, and future perspectives in Section 8.
2 FUNDAMENTALS
This section provides an overview of the fundamental aspects of textiles, detailing the definitions and characterizations of textile objects, exploring the variations in their physical properties, and examining the diversity in manipulation tasks. Understanding these foundational elements is essential for delineating the types of variations and challenges discussed in the subsequent sections on modeling, perception, and manipulation.
2.1 Textiles, Fabrics, and Cloths
Figure 1: Manufacturing process of textiles: textile is an umbrella term that covers materials that are made of interlacing natural or synthetic fibers. The figure depicts the textile production process in its different stages, where yellow boxes represent materials and objects, while blue blocks specify the processing step.
{marginnote}
[…]
Textile has evolved from its initial reference, woven fabrics, to encompass a broad spectrum of objects. These include traditional woven fabrics, as well as deformable and flexible materials made from yarns or threads through various construction processes beyond weaving. Common terms used to describe textile objects are fabric and cloth. Despite being often overlapped in usage, they carry subtle distinctions tied to different phases of the production process, as shown in Fig. 1. Fabric is any thin, flexible material crafted from thread or yarn, fibers, polymeric film, foam, or their combinations, and is used in creating further products like clothing, requiring additional production steps [15]. Cloth, while sometimes used interchangeably with fabric, typically refers to fabric that has undergone further processing or cutting. Everyday clothing items are predominantly made through weaving or knitting before being sewn, with high fashion exploring other methods more extensively.
The physical and mechanical properties of the final textile object are intrinsically tied to its manufacturing process. Yarns and threads, the building blocks of these textiles, are spun from various origins—animal, plant, mineral, synthetic, or blends thereof. Woven fabrics, produced by interlacing two sets of threads, are known for their hardness and non-elasticity, making them ideal for garments like shirts and jeans that retain creases. In contrast, knitted fabrics, created by interloping a single set of yarn, offer softness and elasticity in all directions and thus are used in wrinkle-resistant clothing such as t-shirts, which do not stretch uniformly in all directions to better conform to the body’s shape. Additionally, further production steps involve cutting and coloring fabrics to create clothing with various shapes and colored textures. A comprehensive exploration of fabric properties is given in [16].
2.2 Variations of Object Properties
One of the major challenges in robotic manipulation is designing planning and control algorithms that generalize or adapt to variations that will be encountered in real-world deployments [17]. In this review, we refer to generalization as the ability to apply the knowledge acquired from a specific set of textile properties to unseen variations of these properties. Adaptation is instead the ability to dynamically adjust strategies or models in response to changing conditions or novel variations in the object properties. Given the challenges of generalization and adaptation, it is crucial to first understand the variations in textile properties, which can be broadly classified into physical and mechanical properties. Fig. 2 shows an example of different variations of properties of textiles for four different objects.
2.2.1 Physical properties
The object’s physical properties are inherent characteristics and can be observed and measured without subjecting the material to external forces or manipulations. These properties describe how a material appears under static or non-changing conditions. Among these properties, Size and Shape are crucial as they influence the robot’s workspace, correlating with the object’s function and category. The Weight influences how the textile deforms under gravity. For heavy and large textiles such as blankets, robots may need bimanual systems to manipulate them. The Color affects perception-based algorithms, essential for identifying and tracking the textile during manipulation tasks [1]. Fabric Material composition, ranging from natural to synthetic fibers, dictates interaction forces between the textile and the robot, altering manipulation strategies [18]. The Construction Technique, referring to the knitting or weaving processes, directly impacts a fabric’s mechanical behavior by defining, for example, how thigh or loose the construction pattern is, thus influencing the object elasticity and rigidity [19].
2.2.2 Mechanical properties
The mechanical properties are parameters that describe how a material responds to applied forces or manipulations. Cloth Stiffness or rigidity influences how it behaves under manipulation as it determines the resistance to deformation. Stiffness is a key determinant of a fabric’s manipulation behavior, as it quantifies the textile’s resistance to bending and deformation. Elasticity, or the capacity of a textile to stretch and recover to its original size after being deformed, is essential, particularly in applications like robotic dressing assistance [20]. High elasticity can accommodate user movements and minimize safety risks during interaction. Elasticity and stiffness are properties of the material that are often characterized by Young’s modulus and Poisson’s ratio. The Young’s modulus quantifies a material’s ability to resist deformation under stress by showing the relation between stress and strain in the elastic region of the textile. The Poisson’s ratio measures instead the ratio of transverse strain to axial strain in a stretched material. Finally, Friction refers to the resistance encountered when one surface slides over another. In the context of textiles, friction can vary significantly depending on the surface characteristics of the fabric [21]. Friction properties affect how textiles move against surfaces, thus impacting how robots grasp and transport these materials and how they interact in contact with human skin or other clothing layers.
These mechanical properties vary based on the production process the textile undergoes. They can be measured through standardized methods from the textile industry [22] or through procedures tailored to robotics applications [23]. Their interdependent influence in cloth deformation poses fundamental challenges to the analysis and understanding of how these properties affect robotic manipulations.
Figure 2: Object properties variations: Object differences rely on the variation of their physical and mechanical properties. The figure shows visual examples of how properties can vary between four objects from literature cloth sets. Measures can be found in [23].
2.3 Variations of Tasks
Manipulation of textiles spans a wide range of applications across domestic, healthcare, and industrial settings, including tasks such as laundry, tidying, dressing, sorting, and more. Each of these tasks presents unique challenges that necessitate the design and study of specific manipulation strategies that account for specific variations in the physical and mechanical properties of textiles discussed in the previous sections.
Following the taxonomy proposed by Mason [24], manipulation techniques can be characterized based on the nature of forces involved: kinematics, static forces, quasi-static forces, and acceleration forces. In the context of manipulating deformable objects, most current techniques can be broadly classified into two main categories [25]: quasi-static manipulation, which involves slow motions allowing static equilibrium, and dynamic manipulation, which involves motions that include acceleration forces. Physical properties like material, as well as shape and size, significantly affect both types of manipulations. On the other hand, while mechanical properties play a fundamental role in tasks requiring dynamic manipulations due to the influence of acceleration forces, tasks involving quasi-static manipulations are less dependent on mechanical properties.
For instance, tasks such as classification or sorting highly depend on physical properties such as the shape of the cloth, as well as the material the textile is composed of [26]. Similarly, tasks like flattening and folding rely heavily on understanding and manipulating the shape of the textile [27, 28, 29, 30]. In these cases, elasticity does not directly impact the outcome of the manipulation. However, stiffness and friction still play a role in the deformed state in which the object is observed. Regarding shape properties, flattening and folding tasks are typically performed on textiles laid flat on a surface, simplifying the manipulation process by reducing the need to consider more complex topographical features such as loops, holes, or cylindrical parts. In contrast, tasks such as assistive dressing [31] or hanging an apron on a hook [32] require a nuanced understanding and control of the textile’s topological features, including connectivity, holes, voids, and spatial relationships [33]. These tasks are inherently more complex due to the need to manipulate the textile in three-dimensional space and adapt to its changing configuration. Mechanical properties, such as elasticity and stiffness, play a crucial role in these scenarios, as they influence how the textile drapes, stretches and recovers from deformation.
3 MODELING
Accurate modeling physical properties of textile objects is crucial for robotic manipulation, computer graphics, and material sciences. In this section, we introduce modeling techniques commonly used in robotics, focusing on how they characterize various properties and their effectiveness in capturing them. Rather than elaborating on the technical aspects of these techniques, which are extensively covered in other reviews, we aim to provide an overview of their utility and the types of physical and mechanical variations they address. The reality gap within commonly used simulators is also covered with a discussion about currently unmodeled phenomena and techniques for real-world alignment. For an in-depth technical discussion on textile modeling techniques, we refer readers to [14, 13, 34].
3.1 Physics-based Models
Physic-based models form the foundation for the most commonly used simulators in robotics. These models usually describe the cloth state through geometric representations such as particles or meshes. Physics-based simulation provides a reproducible and scalable mechanism to study physics-based cloth models and cloth dynamics across a range of environments and tasks. In what follows, we characterize different modeling techniques based on the representation used for the cloth.
Mesh-based models, such as mass-spring-damper models, represent the cloth as a mesh of interconnected masses and springs. The physical interaction between masses intuitively reflects the variation of textile stiffness and elasticity. These models are a popular choice implemented in many simulators, for example, MuJoCo [35] and SOFA [36]. While these models are rather straightforward to implement, they are most suitable for simulating small deformation but not complex elastic effects with high fidelity. Additionally, tuning parameters for the springs and dampers to achieve realistic behavior can be relatively challenging, as these parameters do not directly correspond to physical meaning in the real cloth.
Particle-based methods in cloth simulation represent the material as a collection of discrete particles, each defined by properties like position, velocity, and mass. These particles are interconnected through holonomic constraints that describe their interactions. A widely used approach within this framework is Position-Based Dynamics (PBD), as implemented in simulation engines such as FLEX [37, 38]. PBD focuses on satisfying constraints related to stiffness, elasticity, and collision by directly adjusting the positions of particles, with subsequent velocity updates derived from these positional changes. The advantages of PBD include rapid simulation speeds, enhanced stability, and the ability to simulate inextensible textiles and shear deformation. The ability to easily incorporate a variety of constraints makes it possible to model plastic deformations, as well as interactions with fluids and rigid bodies under flexible frictional models [37]. While PBD offers visually plausible dynamics, it may not always align with physical realism. Moreover, the interpretation of simulation parameters into physical parameters, such as material modulus, remains challenging and often necessitates extensive tuning to achieve desired effects.
Continuous domain models leverage the principles of continuum mechanics rather than relying on discrete representations of particles or meshes. This approach treats physical quantities of objects as continuous fields, providing a more physically precise depiction of material deformations. The underlying partial-differential equations are commonly solved with the finite element method, which divides the field domain into small elements linked as meshes. While this approach can be computationally intensive, model parameters have a clearer physical interpretation for textile properties such as strain, stress and their relation through Young’s modulus. Robotics-relevant simulators relying on this modeling technique are SOFA [36], Isaac Sim [39], and Bullet [40] but specifically for deformable objects. Recently, the Material Point Method (MPM) [41] emerges with a mix of particle representation, for which the particles carry material properties, while a spatial grid is used to compute forces and update the state of particles. By combining both representations, MPM is promising for rapid simulation of complex deformations observed in a cloth.
Besides forward prediction, differentiable simulators have recently gained popularity with efficient derivative evaluation through automatic-differentiation or adjoint methods [42]. The feature extends existing model-based approaches for efficient policy learning and parameter identification, finding applications in cloth-related tasks [43, 44].
3.2 Data-Driven Models
Data-driven models have become a prominent alternative to traditional physics-based methods for simulating complex materials and deformations. Compared to physics-based techniques, data-driven models offer greater flexibility in defining the state space for modeling cloth dynamics, as well as lower computational complexity and easier parallelization.
These models learn cloth dynamics directly from data, which can range from handcrafted observations, such as cloth key points, to raw images and 3D point clouds or a latent representation of these. Image-based inputs have the benefit of not requiring knowledge of the 3D state of the cloth as the dynamics are learned either in the pixel space [30] or in a latent representation of the image [4]. It also offers the advantage of working under partial observability [30]. However, these often struggle with domain shifts that can occur with changes in camera positions, lighting, or background conditions, which impacts their generalization to new environments. Particle-based representations are often more robust to changes in visual conditions as they rely on 3D geometric representations such as particles or meshes rather than visual clues. These representations require specific architectures to efficiently capture local structures and handle data sparsity. Examples include PointNet++ [45] for unordered point sets and Graph Neural Networks (GNNs) [46] for mesh-based representations. Despite the advantage of particle-based representation in generalization, these models are generally more computationally intensive than their image-based counterparts. In addition, registering a mesh from partial depth data could also be non-trivial for textiles under large deformation.
The flexibility of data-driven models allows one to address variations of cloth properties by designing appropriate training schemes or conditioning the models on the object properties or a latent representation of these. Invariance to color changes is typically achieved during augmented training with domain randomization [47]. In contrast, mechanical properties like stiffness and elasticity are commonly addressed through explicit or implicit conditioning. Explicit conditioning estimates these properties or their latent representations through perception and concatenates the estimation to the input model [19]. Specifically for graph or mesh representations, properties like stiffness and elasticity can be induced as edge features in the mesh structure to bias model learning. Implicit conditioning, on the other hand, incorporates these properties by conditioning the model on a recent history of observations or employing techniques such as intuitive physics [48]. These approaches allow models to adapt to changes in mechanical properties based on observed behaviors.
3.3 Reality Gap in Cloth Models
Different modeling techniques often need to account for a trade-off between computational cost and model accuracy. Trading accuracy for speed often leads to a gap between the dynamics in simulations and those in real-world scenarios. This gap between the model and how the physical cloth behaves in the real world is particularly relevant to real-world textile manipulation. Here, we identify gaps between the assumptions made by commonly used robot simulation and real textile properties, as well as progress to align simulation to the real world.
3.3.1 Unmodeled Phenomenas
Robot simulators generally favour fast and visually correct behaviours at the expense of physical realism. Geometrically, textiles are often modelled as 2D objects without thickness. This requires extra care when the task concerns stacking or folding textiles of multiple layers. The mechanical modelling, be it physics-based or data-driven, rarely goes down to the basic building blocks of fabrics as laid out in Section 2.1. The simulation hence neglects aspects like fabric weave, thread count, as well as the response to real-world factors such as humidity, temperature, or wear and tear. In computer graphics, yarn-level simulation [49, 50] has been studied to account for warp and weft interactions. However, this has not yet been integrated into robotics modeling pipelines. The simplification of simulating effects in smaller granularity also impacts the fidelity of how a cloth interacts with external objects. Textiles likely exhibit more complex sliding behaviours along different directions due to how they are constructed. This breaks the popular isotropic friction cone model in many simulators. Missing yarn-level simulation also influences modeling the tear-up of a cloth. Current robot simulators often assume invariant topology and uniformity about how a cloth is constructed. Graphics research has demonstrated the possibility of capturing these nuances with advanced simulation [51, 49].
3.3.2 Aligning Simulation to the Real World
Physics simulation is essential for generating synthetic training data, exploring learned policies, and predicting the performance before real-world deployment. Thus, bridging the gap between the real world and simulation is crucial for generalizing manipulation skills to real-world textiles. A common approach to reduce this gap involves tuning the simulation parameters to align with reality. This process is often referred to as real2sim, or system identification, and can be classified into three categories [13]: gradient-based techniques, global optimization techniques, and neural network-based techniques. For gradient-based optimization, differentiable simulators backpropagate the error between the real-cloth state and the simulated one for system identification. These methods, however, often face challenges with discontinuous loss landscapes, particularly in scenarios involving deformable objects. For scenarios where the analytical model is not differentiable, global optimization techniques such as Bayesian Optimization (BO) and Covariance Matrix Adaptation Evolution Strategies (CMA-ES) are often used without the need for gradients [52, 53]. Global search methods can be computationally intensive and may struggle with scalability in high-dimensional spaces. A combination of the two has been proposed in [54], where they combine global search using BO with a semi-local search to retain the benefit of gradient-based optimization but integrate BO for the parts of the landscape that are intractable for gradient descent alone. Another approach is to combine Bayesian inference with neural networks to infer simulation parameters [55]. One of the major benefits of this class of methods is capturing uncertainty in parameter estimates.
4 PROPERTIES PERCEPTION
The perceptual capabilities of robots encompass a variety of skills, including state estimation, segmentation, tracking, recognition, classification, and the identification of appropriate grasping points on cloth items. A comprehensive review of perception for grasping is presented in [11]. Additionally, discussions on state estimation, parameter identification, and detection are provided in [14], where the focus was not only on textiles but on deformable objects in general. This section will specifically delve into the identification of textile properties, aiming to focus on perceptual capabilities that enable the generalization and adaptability of robots to variations in the textile properties introduced in prior sections.
Although robots in human environments constantly face uncertainty about object properties due to changing conditions, perceptual systems can reduce this uncertainty by providing sensory feedback [56]. Perception can be passive, requiring no physical interaction with the objects of interest. Properties such as color, shape, and material can be estimated using passive perception [57]. However, properties such as elasticity and friction cannot be observed in static scenarios. Therefore, interactive perception is necessary to obtain accurate estimates of these properties  [58].
Designing the perception process relies on several key factors: the relevant properties for the tasks at hand, the available sensors, and the manipulator configuration. Figure 3 provides an overview of various design choices for estimating textile properties. This table also represents the literature discussed in the remainder of the section.
Figure 3: Overview of sensors, properties, and actions in passive and interactive perception for textile manipulation. On the left, the figure presents a list of sensors, properties, and actions. The different colors of the sensors reflect to what extent the sensor has been explored for textile perception. On the right, it establishes connections between these categories based on the current literature, demonstrating how passive perception methods, using sensors like cameras and spectrometers, estimate properties that do not require physical interaction. In contrast, interactive perception integrates actions to improve the robot’s understanding of mechanical properties, leveraging specific actions coupled with sensor feedback to accurately perceive properties such as weight, stiffness, and elasticity.
4.1 Physical Properties Perception
As noted in Section 2.2.1, physical properties like size, shape, weight, color, fabric material, and construction technique can often be observed and measured without external manipulation. Particularly, size and shape can be effectively inferred by observing textiles spread over a flat surface and utilizing contour or segmentation techniques [59]. Data-driven techniques such as land-mark detection or end-to-end classifiers have also been explored for shape classification [60, 61]. However, textiles frequently arrive in deformed or crumpled states, complicating direct visual assessment. While some approaches have explored feature extraction processes to classify clothes from highly crumpled configurations [26], interactive perception is often necessary to enhance perceptual accuracy [62]. Lifting interactions, for example, are commonly employed to infer the shape of the textile better and facilitate its classification by garment type  [63, 64].
The estimation of cloth weight has been relatively underexplored within the robotics community. Pioneering investigations, primarily from the computer graphics domain, have demonstrated the potential of estimating the weight per area of textiles through video analysis of the object being influenced by external forces like wind [65, 66]. This interaction could be analogously implemented in robotic systems as a flinging action [67], although this specific application has yet to be explored for weight estimation in a robotic context. Current approaches in the robotic community estimate cloths weights through the estimation of external forces/torques felt by the robot during motion tasks such as lifting [68]. An alternative approach is to rely on a similarity network to infer cloth weight from lifting interactions using ground-truth cloth weight as supervision [62].
Textile material and construction techniques have received considerable attention in the robotic community to explore combinations of visual and haptic sensing specifically [69]. Photometric stereo sensors can reconstruct the surface of a region of the cloth at the yarn level and exploit the 3D pattern to infer the material and construction properties of the object [70]. Alternatively, high-resolution tactile sensing such as GelSight [71] has shown to be relevant for both material classification and construction technique classification of fabrics. These sensors provide high-resolution images, by pressing the sensor on the cloth [72]. More specialized robot sensors, such as micro spectrometers, offer detailed insights into the yarn material composing the fabric [73].
Since both material and construction techniques influence the mechanical properties of cloth, interactive perception can provide additional insights to identify these properties accurately. Low-resolution tactile sensors like the BioTac, coupled with a contact microphone, have been explored to infer the material of textiles by exploring the feature of the signal recorded from a sliding interaction over the textile [69]. Additionally, force-torque sensors used in conjunction with pulling and twisting actions have provided further insights into materials and construction techniques by observing how the textile responds to different mechanical stresses [18].
4.2 Mechanical Properties Perception
Mechanical properties of textiles, such as stiffness and elasticity, are critical in determining how a fabric will behave under various types of stress and strain. These properties are traditionally measured using sophisticated and costly systems that are standard in the textile industry. A prominent example is the Kawabata Evaluation System for Fabrics, which quantifies a fabric’s response to controlled forces by measuring parameters such as stress-strain behavior at maximum load for the specific material being tested [74]. Additionally, the computer graphics community has conducted in-depth studies of cloth elasticity using setups that combine pulling interactions with force and visual observations [75, 76]. However, these setups are typically unsuitable for robotic applications due to their lack of real-time interaction and computational capabilities.
In the robotics community, stiffness and elasticity have been explored using a combination of force-torque sensors and camera observations during robotic pulling interactions [19]. While this approach directly measures how textiles respond to controlled mechanical stresses, significant research focuses on methods using only camera observations to detect textile deformation under external forces like gravity or wind. Specifically, work using cameras can be broadly divided into two categories: 1) optimizing simulation parameters to reflect the behavior of real-world objects accurately and 2) applying external forces and observing resultant shape changes. The first category leverages 3D observation, such as point clouds, to optimize simulation parameters such as elasticity, bending, and stiffness [77, 78]. The second category instead includes methods from the computer vision community, where 2D images are used to assess how a textile drapes under gravity or moves under external forces. These observations focus on visual features such as wrinkles or the dynamic motion of textiles over time [42, 79, 80]. Notably, the studies of this last category do not involve robots, indicating a potential new research direction to replicate these techniques in robotic systems. Friction, on the other hand, remains largely underexplored. While some work has investigated haptic adjectives related to friction, direct estimation techniques are limited, with efforts mainly focusing on using low-resolution haptic and auditory sensors to infer friction-related features [81].
5 MANIPULATION
Robotic manipulation in human environments often confronts highly unstructured settings, necessitating the ability to manipulate objects with varied characteristics. In Section 2, we reviewed textile variations in terms of properties and tasks, and highlighted the complexity of addressing the variations of these real-world textiles. These complex environments underscore the importance of endowing robots with robust adaptation capabilities. Here, adaptation refers to the ability of the robot to adjust its strategies in response to changes in its operating environment or variations in the objects it manipulates. Effective manipulation skills require adaptability to both previously observed and novel variations of textile properties [56].
Successfully handling different clothing variations can be achieved through two primary mechanisms: generalization and adaptation [17]. Generalization assumes that the distribution of knowledge acquired in the source domain will encompass the distribution of the target domain, resulting in a successful transfer. This assumption allows the robot to handle known variations, but it may not hold in the presence of novel variations typical of diverse human environments, where previously learned knowledge may fall short. A common framework to achieve generalization is domain randomization [47]. In contrast, adaptation involves dynamically modifying the manipulation strategy or dynamics model by taking into account environmental changes. This adaptation can be achieved by parameterizing the dynamics model [82, 48] or the manipulation policy with the estimated cloth properties.
In the following sections, we group common manipulation strategies into three categories, provide a high-level overview of each, and discuss in more detail their adaptability towards different textile properties and task variations. A general overview of these methods is given in Table 1, including the tasks these are applied to as well as the variations of cloth physical and mechanical properties that are evaluated.
Table 1: Overview of textile variations handled by methods in the literature. Methods that do not assess their generalization or adaptation over textile variations are marked with , methods that address textile variations explicitly with , and with () when not explicitly mentioned but the evaluated textiles contains such variations.
Ref. # O Task Physical Properties Mechanical Properties
SH SZ CLR M CT W E ST F
Model-Based
[4] 1 Fo
[9] 1 Fo/Fl
[48] 6 Fo
[21] 7 Fo () ()
[30] 1 Fl
[83] 3 Fl
[84] 3 Fl () () () () ()
[85] 5 Fl () () () () ()
[53] 1 D
Model-Free
[7, 28, 8] 1 Fo
[86] 2 Fo
[87] 3 Fo
[88] 3 Fo () () ()
[89] 3 Fo
[90] 5 Fo ()
[91] 10 Fo () () ()
[92] 10 Fo/Fl ()
[93] 20 Fo () () () () ()
[94] 1 Fl
[47] 3 Fl
[67] 3 Fl
[95] 3 Fl
[96, 97] 1 D
[20] 2 D () () () () ()
[98] 3 D
[31] 5 D () () () () ()
[99] 3 B ()
Heuristic-Based
[100] 4 Fo ()
[29] 12 Fo () () () () ()
[101] 25 Fo () () () ()
5.1 Model-based Manipulation
Model-based manipulation approaches refer to the class of methods that use a model of the textile to generate the manipulation strategy. The manipulation actions are usually obtained by planning with the model [53, 84, 85]. These approaches can be categorized based on the methods used to construct the model, namely into analytical models and data-driven models.
5.1.1 Analytic Models
Analytic models, or physics-based models, build the textile model based on the laws of physics. Please refer to Section 3.1 for a more detailed discussion on the popular analytic models and the textile properties they can model. When using these models for manipulating a textile object, the initial step typically involves aligning the model dynamics to the real world to match the dynamics of the target textile [21, 102]. Analytic models have the advantage of being applicable to different tasks as the model is usually task-agnostic, as long as the textile dynamics required by the task can be accurately modelled by the chosen analytic model.
There are two major drawbacks of using analytic models for manipulation. First, their generalization is limited by the accuracy of the physics model, which, as discussed in Section 3, is often only an approximation of real cloth dynamics [52]. Second, each new object requires a separate system identification process, which can be tedious and difficult to scale for adapting to a great variability of textile mechanical properties.
5.1.2 Data-driven Models
Data-driven models learn from data of interactions with the textile. There can be many state representations for data-driven models as laid out in Section 3.2. The choice of the state representation and model architecture can greatly affect the generalization ability of the model towards different properties such as the shape and texture of the textile. A particle-based state representation combined with graph neural networks [84, 85] has been shown to generalize better towards different shapes, geometries, and mechanical properties of the cloth compared to images [30] or latent-based state representations [83, 4], as the particle-based state representations align better with the underlying cloth physics. Using depth images [9] or point clouds [84, 103, 85] as the state representation also naturally makes the model invariant to the color and visual texture of the textiles. Due to the large amount of data needed for learning the model, most works use a simulator to generate the interaction data for learning the model [30, 83, 9, 84, 48, 85], with a few that does so in the real world [4].
As discussed in Section 3.2, data-driven models often address variations of cloth physical properties through domain randomization [84, 85, 83, 30]. However, as current simulators [35, 40, 37] struggle in modeling a wide range of mechanical properties of the textiles, the generalization towards diverse mechanical properties that can be achieved through domain randomization remains limited. Still, recent work has shown that integrating a model conditioned on a latent representation of mechanical properties within a feedback-loop framework enables a model learned in simulation to adapt to textiles with diverse mechanical properties in the real world [48].
Most works in the literature learn an individual model for each type of manipulation task, e.g., assistive dressing [53], cloth folding [48, 9], cloth smoothing [9, 83, 85] or blanket covering [103]. A few works have demonstrated that the learned model can be used to perform two tasks such as cloth folding and smoothing [30, 84, 9]. Learning a single model that can generalize to multiple manipulation tasks remains underexplored.
5.2 Model-free Manipulation
Model-free manipulation approaches directly map the state or sensory observation of the textile to the manipulation action without having a model in the loop during inference. Compared to model-based approaches, they require no prior knowledge of the textile as no model needs to be constructed. There are two main approaches: reinforcement learning and imitation learning.
5.2.1 Reinforcement Learning
To apply reinforcement learning (RL) to a textile manipulation problem, the manipulation problem needs to be formulated as a Markov Decision Process (MDP) or a Partially Observable MDP (POMDP). The key in the formulation is the design of the state, action, and reward space of the MDP. As in the case for model-based methods, common choices for state representations of textiles include manually defined features such as key points on the cloth [91, 96], or results from a perception system such as an image [47, 7, 88] or point cloud [20, 31] of the textile. Common actions for manipulating textiles include action primitives such as pick-and-placing [47], dragging [89], flinging [67], or raw actions that control the delta movement or velocity of the robot end-effector [88, 7, 31]. The reward function defines the desired outcome of the manipulation task based on the state of the environment.
Given the formulated MDP, reinforcement learning algorithms can be used for finding a policy that maximizes the expected accumulated rewards, thus learning the necessary manipulation skills [47, 7, 31, 88, 95]. Most works formulate the textile manipulation problem as a multi-step MDP [47, 31, 7, 88, 95] and learn a policy to maximize the accumulated reward, while some works [67, 89] formulate the problem into a bandit problem (an MDP with only 1 step), learn the reward function (usually in the form of a spatial-action value map [67, 28, 89]), and choose the best action as the one that maximizes the learned 1-step reward. Due to the large amount of data required by an RL algorithm, the need for a reward function, and the need to periodically reset the environment, such methods usually train a policy in a simulator and perform sim2real transfer [47, 7, 31, 91, 88, 95] with optional real-world fine-tuning [67]. A few works [86, 89, 28] directly train in the real world, where the reward can be automatically computed from real-world perception systems [89], and the environment can be automatically [28, 89], or manually [86] reset.
Domain randomization is still the key technique for model-free RL methods to achieve generalization towards diverse textile mechanical and physical properties. In domain randomization, properties of the textile and environment are randomized during the training process, so the resultant policy generalizes to all varied properties. The randomized quantities can include the textile’s shape [31, 67], size [31, 7], location and orientation [31, 7], texture and lighting [7, 88, 47], and mechanical properties [88, 47, 91]. To achieve more informed domain randomization, contrastive learning can be used to compare pairs of real and simulated garment observations to learn a similarity metric [97], which is used to tune the simulation parameters to align the simulation and real-world garment observations. Again, one caveat of domain randomization is that the range of mechanical properties that can be randomized is limited by the fidelity of the simulator. There has been little work in the literature that explores addressing variations of textile properties via adaptation.
In terms of tasks, model-free RL methods have been applied to many textile manipulation problems including cloth smoothing and flattening [38, 47, 67, 89, 95], folding [86, 7, 89, 28, 88], placing [38], hanging [7, 32], blanket covering [104], and assistive dressing [20, 97, 31, 96]. All these works learn an individual policy for each task; the goal of learning a single policy over multiple tasks or that can generalize to different manipulation tasks remains highly unexplored.
5.2.2 Imitation Learning
To apply imitation learning (IL) to textile manipulation problems, a dataset of expert demonstrations that solve the manipulation task needs to be first collected. Such demonstrations can be collected by a human [93, 90, 98] or using a scripted policy [94, 105, 87, 3]. The most standard IL algorithm is behavioral cloning  [93, 90, 94]. Another approach is to directly map the actions in the demonstrations to the test object via learned correspondences [92].
The generalization ability of IL approaches highly depends on the amount of variations presented in the demonstrations. An extensive dataset that covers diverse objects with varied properties is essential for learning a policy that can generalize across different textiles. Many works that use IL for cloth manipulation show limited generalization towards different textile properties, demonstrating their method on a fixed textile [105, 94, 8, 3]. Some works use depth images as observations, so the policy can be invariant to visual features such as color and texture [90, 99, 87]. When the collected demonstrations are diverse enough, a stronger generalization can be achieved: Xue et al. [93] collected a dataset with hundreds of simulation shirts and 40 real-world shirts, and showed that the imitation policy can generalize to 20 real-world shirts with diverse shapes and materials. One caveat of IL methods is that collecting a sufficiently diverse and extensive set of demonstrations can be human labor-intensive.
Task-wise, IL methods have been applied to cloth smoothing [94, 3, 92], folding [8, 93, 90, 87, 92], twisting [105], bed-making [99], and dressing [98]. Again, most works have only learned a single policy for one task [94, 8, 105, 90], with only a few that either learns a single backbone and different output branches on the same backbone [93], or a shared correspondence [92] for different tasks.
5.3 Heuristic-based Manipulation
The last discussed approach in textile manipulation is based on human-designed heuristic rules instead of learning the manipulation strategies from data. The heuristic rules vary depending on the target manipulation task. For example, for cloth smoothing, one commonly used rule is to detect the wrinkles of the cloth, and the manipulation action pulls the cloth in the perpendicular direction to the detected wrinkle direction [99]. For grasping, one rule is to use the position and orientation of a wrinkle to compute the target location and orientation of the gripper [29]. Another heuristic for a cloth article hanging in the air is to grasp key points that are identified based on the border geometry, e.g., corners for a towel, or sleeves for a t-shirt [101]. For folding, a heuristic folding motion can be achieved by leveraging gravity and moving the cloth such that the moved part is always vertical, which is called a “g-fold” [100]. For assistive dressing, common heuristic solution is to move the grasped garment along the forward direction of the human limb [53].
The generalization ability of such heuristically defined rules vary based on the target manipulation task and the assumptions made. Usually, these methods generalize well under the settings where the assumptions are satisfied, and tend to have limited generalization when the assumptions are broken. Due to the diverse properties and configurations a textile can have, it can be hard to design a heuristic rule that would generalize to every situation one might encounter when manipulating textiles.
6 BENCHMARKS AND DATASETS
Benchmarking robotic manipulation plays a crucial role in understanding methods’ capabilities and limitations enabling standardized comparison. Particularly benchmarking of cloth manipulation is limited due to the absence of objective and consistent evaluation processes and the limited research on how different textile materials, features and structures influence the performance of a method.
In the following subsections, we will discuss the two main tools for benchmarking: datasets and benchmarks. Datasets can demonstrate the adaptability of a method to new data. However, when physical interaction is inherent to the task, the generalization of a method needs to be demonstrated through a benchmarked experimental validation, including standardization of the objects used.
6.1 Benchmarks
Most benchmarks tackling cloth objects evaluate manipulation in simulated environments. Simulation benchmarks have the advantage that they can automatically generate many different conditions, enabling the evaluation of generalization. For instance, SoftGym [38] is a benchmark for deep reinforcement learning of deformable object manipulation that provides a set of simulated standardized environments and includes a set of cloth manipulation tasks. Assistive Gym [106] is a physics simulation framework for assistive tasks that includes an assistive dressing task with a hospital gown. However, due to the sim2real gap, it is necessary to have benchmarks for evaluating real-world applications with physical objects. Garcia-Camacho et al. [107] proposes benchmarks with real executions for three cloth manipulation tasks including protocols, qualitative evaluation metrics, and several complexity levels for each task based on the initial state of the cloth. Alternatively, Clark et al. [108] proposes four benchmark tasks for evaluating the performance of end-effectors in grasping clothing items, along with protocols to normalize starting crumpled configurations and metrics. Both works agree that an important aspect of benchmarks is the standardization of the objects. In addition, an effective way to measure the generalization of a method is through the use of a wide number of objects with varied properties. As it was done for rigid objects with the now widespread YCB object set [109], some efforts were put into replicating it for clothes [110], providing an extension of the YCB object set focusing on textile objects. It includes a wide variety of cloth household objects with benchmarking guidelines for its use. This object set was distributed among research institutes and the participants of the cloth manipulation and perception competition organized in IROS’22 and ICRA’23 to extend its usage, a requirement for a benchmark to be useful. However, there is an important issue in defining standardized textile object sets given by stock continuity since the fast-fashion industry changes stock each season, preventing the maintenance of the same objects for extended time periods. To solve this issue, a method for building comparable textile object sets across different publications has been proposed in [23]. It proposes a framework to characterize textile objects through the measurement of their physical and mechanical properties, enabling the quantification of variability on the textile properties listed in Section 2 of a given set of objects. With this characterization, the generalization of a method can be quantified based on the amount of variation that the objects offer. This idea goes in line with the one proposed for rigid objects in [111], where generalization is measured similarly but also adds other aspects of variability in background, table color, etc. Variability in the starting configurations and scene properties of color and light should also be considered for cloth manipulation.
6.2 Datasets
Unlike benchmarks, datasets serve the purpose of providing data for designing or learning a task. The extent of generalization depends on the variation covered by the dataset. Datasets are often task-specific, ensuring that the data is relevant both for training a system and evaluating its performance. In cloth manipulation, the most common datasets include simulated 3D models [112, 113], RGB images [61, 114, 115] or depth [89].
The largest existing datasets are for classic vision problems like cloth classification and landmark point detection, with datasets such as ClothesNet [112] in simulation or DeepFashion [61] with real images and segmentation of people wearing all type of clothes with datasets such as CLOTH3D [113] in simulation or [116] with real images. Surface reconstruction is another classic perception problem that has been applied to cloth [114]. For cloth classification or landmark detection, images contain annotations of cloth type and location of landmarks, but usually, images come from the fashion industry, and so clothes are either flatted, hung on hangers, or worn by humans. For surface reconstruction, annotations need to have realistic images with the real mesh of the object, and therefore, existing datasets are all in simulation and subject to the sim2real gap discussed in Section 3.3.
For robotics, cloth classification and landmark point detection are also important, but they need to be identified during the stages of manipulation where clothes are in very complex configurations. Indeed, cloth classification is required but from crumpled states [26], or when grasped by one point [117]. Other features that need to be identified are corners and edges [118] or wrinkles [119]. The dataset [115] is an expansion from DeepFashion [61] to adapt it to robotic manipulation.
One of the challenges for datasets in cloth manipulation is labelling the ground truth of the deformation state in real images. Only a few datasets exist with point clouds with labels of what points correspond to corners or edges during a manipulation [120]. So far, to encode interactions, datasets use RGB-D images where the action is annotated as a pick-up pixel point and a direction of motion in the image [89], with the limitation of only representing close-to-planar configurations. More complex actions appeared lately with Visual Language Models where a sequence of images is linked to a sequence of positions of the end-effector [121], with corresponding datasets.
Datasets rarely annotate the variability in the mechanical properties mentioned in Section 2, and some of the physical properties are covered depending on the requirements of the trained system. The main issue is the difficulty in labeling the deformation ground truth from images due to the severe self-occlusions, while datasets in simulation are only partly useful due to the sim2real gap.
7 APPLICATIONS AREAS
We outline the applications and challenges of manipulating deformable textile objects in scenarios requiring robots to adapt and generalize to their varying properties. Understanding the physical properties of textiles is crucial for a wide range of tasks in diverse sectors such as household chores, healthcare, and the fashion industry. Table 2 provides a (non-exhaustive) overview of work categories in these sectors, organized by the frequency at which they are addressed in the literature. Tasks such as folding, smoothing, and dressing receive frequent attention from the community, whereas tasks like buttoning, dyeing, and washing remain rather underexplored. In what follows, we will discuss in detail these tasks and the requirements concerning variations of textile properties.
Table 2: Overview of Variation of Tasks addressed by the community and their frequency.
Frequency Household Healthcare Fashion Industry
Frequent (4+) Folding [87, 29, 122, 89, 8, 92, 123]
Smoothing [67, 92, 30, 99, 47]
Ironing [124, 125, 126, 127, 128]
Dressing [123, 129, 123, 130, 131, 132, 53, 133, 134, 31]
Rare(2-3) Hanging [33, 7]
Sorting [26, 70]
Wiping [135, 136]
Bedding [137, 122, 104]
Bed-making [137, 122, 99]
Bandaging [138, 82]
Recycling [70, 139]
Unaddressed (0-1) Storing Buttoning [140] Manifacturing [141]
Dyeing [142]
Quality control
Coloring
Washing
7.1 Healthcare
As populations age worldwide [143], there is a growing opportunity for robotic systems that provide physical assistance with activities of daily living (ADLs) and healthcare tasks. Different assistive tasks involve manipulating textile objects, such as in robot-assisted dressing [123], bedding [137], bathing [134], and medical care [138]. This section examines significant advancements in healthcare robotics, particularly robot-assisted dressing, bathing and bedding, bed-making, and medical care, highlighting opportunities and challenges in real-world applications.
Robot-assisted dressing is crucial for individuals with upper or lower extremity mobility impairments, requiring careful manipulation of deformable objects such as garments. This task necessitates the ability to adapt to various object variations while ensuring the safety and comfort of the patient. Research in robot-assisted dressing has led to the development of robots that can assist with putting on shirts [129, 123, 130], pants [131], and footwear [132], with challenges including ensuring physical safety, accurately modeling human-robot interactions during garment occlusions [53, 133, 134], and generalizing to different garments [31].
Bed-making and blanket manipulation represent significant opportunities for cloth handling in caregiving involving large, deformable textiles. Research has led to robotic systems capable of grasping and smoothing fitted sheets [137], folding and arranging blankets and towels [89, 122]. Key to effective bed-making is leveraging physical properties like elasticity of bedsheets during robotic manipulation [99], also for tasks like autonomously covering and uncovering a person in bed [104].
In daily medical care, robotics research has introduced advances in handling soft materials such as gauze for bandaging [138] or adult diapers [144]. These tasks, involving physical contact with the human body, underscore the importance of incorporating various sensory modalities and control techniques for effective manipulation of soft materials [82].
7.2 Household chores
Several instrumental activities of daily living (iADLs), such as laundry, cleaning with towels (fabric or paper), and hanging clothes, require dexterous textile manipulation. Cloth-like objects are ubiquitous in unstructured domestic environments and pose significant challenges to fully automating these activities. This section outlines methodologies and challenges in cloth sorting, smoothing, ironing, folding, hanging, and wiping tasks.
Cloth sorting involves categorizing garments and textiles by attributes such as item class [26], fabric type, construction, color, and quality, generally before washing or recycling. Accurate perception and classification of these variations can enhance sorting efficiency, enabling generalization across different textile batches.While recent methods have incorporated material identification through tactile feedback [70], few integrate physical interaction to accurately discern mechanical properties.
Robotic tasks such as smoothing, ironing, and wiping involve manipulating fabrics under varying physical conditions, which are primarily influenced by properties like friction and elasticity. Smoothing [67, 92, 30, 99, 47], typically performed before folding or wiping, necessitates to account for varying friction as fabrics transition from crumpled to flat states. Ironing further requires adjustments for temperature variations that affect the fabric’s physical properties. Current methods mainly focus on ironing individual garments [124, 125, 126, 127, 128] and often fail to generalize across different fabric types. Wiping involves tackling the challenges posed by variable surface friction, which can be influenced by the presence of dust or liquids on different surfaces [135, 136]. Each task demands adaptive strategies to cope with the dynamic nature of fabric properties, highlighting the need for methods that can accommodate these variations, as discussed in Section  5.
Significant research in robotic manipulation of cloth has focused on robotic folding with practical applications in healthcare and domestic settings [87, 29, 122, 89, 8, 92, 123]. However, challenges such as variability in garment properties (weight, friction, and shape) affect both quasi-static and dynamic manipulation [52], impacting generalization. Notably, material stiffness, as reflected in bending coefficients, significantly influences each fold, potentially accumulating errors and altering outcomes in the folding process [48].
Cloth hanging involves finding a stable configuration for a garment on a hanger [7] by identifying features like holes and loops [33]. However, the impact of the physical properties of the garment on the deformation of these features remains largely underexplored.
7.3 Fashion Industry
The fashion industry presents numerous opportunities for autonomous robots with adaptive capabilities. This subsection focuses on specific applications such as manufacturing, dyeing, and recycling, directly linked to adaptability in handling, perceiving, and quality management of textiles. Given the extensive previous discussion on sorting, smoothing, ironing, and folding, this section will concentrate on these industry-specific applications.
In cloth manifacturing, robots are increasingly integrated into the cutting and sewing stages of garment production [141], with potential advancements enabling them to detect variations in fabric properties like thickness and stretchability. This ability could allow precisely adjust techniques for each material, significantly reducing waste and enhancing resource efficiency, thus improving sustainability and garment quality.
The automation of dyeing processes significantly enhances sustainability by reducing dye and water usage [142]. Currently, robots in dyeing processes are mainly used for loading and unloading yarn bobbins. Augmenting these robots to recognize variations in garment properties could optimize dye application, tailoring it to the specific needs of each garment.
Finally, recycling is crucial for sustainability in the fashion industry [139]. Similar to sorting, enhanced interactive perception of fabric types, construction methods, and material conditions can improve this process by accurately identifying textiles suitable for recycling [70].
8 DISCUSSION AND FUTURE PERSPECTIVES
This section summarizes the main concepts discussed in this review, further identifying open problems and detailing future research directions and grand challenges to foster the development of perception and manipulation skills that generalize to variations of textile properties and manipulation tasks.
8.1 Synergies
In previous sections, we examined modeling, perception and manipulation separately. However, there is significant interplay among these domains, impacting generalization and adaptation capabilities. Here, we review some challenges faced by these components individually and how their integration can be beneficial for generalization.
Offloading most computational effort to modeling, such as using analytical models, allows for offline optimization of manipulation trajectories. This reduces the burden on control strategies and perception, as optimal strategies can faithfully execute the task in an open loop [21]. While cloth models facilitate transferring to novel tasks [84], real-world applications require perception techniques to adjust simulator parameters for each real-world object as in Section 3.3.2. This involves estimating object properties from sensors or feedback and replanning to address inaccuracies, making perception crucial for generalization [82, 48].
Model-free learning techniques such as RL and IL shift the computational complexity to manipulation by learning end-to-end from raw data [88]. While this allows the reduction of the requirements from the perception module, it has to trade sample efficiency and generalizability, which often results in training a single policy for each task, limiting flexibility in varying environments. However, not all the components of a sequential manipulation task need to be learned end-to-end and offloaded to the policy. Sub-tasks within cloth manipulation can be simplified with perception modules. Examples include flattening based on wrinkle [27] or grasp point detection [64], and folding plans based on corner alignment, which delegate specific tasks to perception, bypassing the need for extensive learning.
Conversely, relying solely on perception and heuristic-based approaches hinders generalizability, as it is challenging to design heuristics for all real-world scenarios. Learning policies can address tasks where heuristic-based algorithms fall short. For example, which point to grasp could be addressed by the perception algorithm, how to reach the grasping point can be resolved with a standard planning algorithm, and how to optimize the manipulation once the cloth is grasped can be learned and performed by an end-to-end policy. The underexplored area of switching between learning and heuristic approaches and determining when to use each presents a promising research direction. This is particularly relevant with the advent of foundation models capable of reasoning about semantics and tasks and adapting to rules and human preferences.
8.2 Open Problems
To discuss the generalization and adaptation of current methods for cloth manipulation, we provided an overview of fundamental textile properties that contribute to the variations of real-world textiles, highlighting the complexity of understanding their role in manipulation due to the intertwined role of each of them. However, the influence of each textile property in robotic manipulation is still underexplored. Moreover, the extent to which explicit identification of each property is required remains unclear. An open avenue for research is identifying a subset of pertinent features for different manipulation task, which can be used to determine the extent to which the textile properties need to be identified.
Adding to the complexity of identifying and defining the textile properties is the dynamic nature of these properties under varying conditions. Aging of the cloth, as well as wet, dirty, or dry conditions, can alter a textile’s behavior, emphasizing the need for continuous sensing and adaptive perception systems in autonomous agents. Moreover, physical attributes like thickness, softness, and durability [72] are typically evaluated alongside semantic descriptors, like smooth, absorbent, hairy, and slippery that humans use to characterize textiles [81]. Particularly, with the advances in Large Language Models (LLMs) there is an emerging potential to bridge these descriptive terms with a physical understanding of textile properties [145].
One key aspect for manipulating a variety of textiles and addressing the variation of properties is perception. While interactive perception with different types of sensors and exploratory actions can enhance adaptability and reduce uncertainty in presence of partial observability, it remains underexplored. Multimodal sensing emerges as a pivotal strategy in this context, integrating various sensory inputs like tactile, visual, and auditory data to provide a more holistic understanding of textile properties. This approach holds significant potential, as it allows robots to leverage multiple temporal information sources, compensating for the limitations of individual sensors.
A notable finding from our review of manipulation techniques is the limited amount of works demonstrating effective generalization to a wide range of deformable objects, with only a few exploring adaptation techniques. Most current efforts utilize domain randomization in simulations to enhance sim2real transfer. While promising, these methods often overlook the potential of various domain adaptation techniques that could allow robots to dynamically adjust their manipulation strategies based on real-time environmental feedback and changes in textile properties. Similarly, the use of state-of-the-art models such as diffusion models, in combination with imitation learning techniques, has predominantly focused on rigid objects due to the complexity of state estimation and high-dimensional state spaces of textiles. This leaves open questions about their applicability to deformable objects and the challenges of deploying these methods without assuming rigidity.
Building on the necessity for improved generalization in manipulation techniques, benchmarking, standardization and datasets emerge as essential yet challenging areas. The diversity of robotic embodiments, sensor configurations, and test sets, complicates the comprehensive comparison of approaches and algorithms. Specifically, the development of universally applicable testing datasets is hindered by issues like stock availability and the difficulties associated with measuring properties accurately. Although there has been some progress towards creating standardized test sets and benchmarks, these continue to be significant open problems within the field. Additionally, a major challenges in dataset availability and scalability is accurately labeling the deformable object dataset, which limits their widespread distribution and application.
8.3 Grand Challenges
To push the boundaries of robotic manipulation of textiles, we identify the following critical challenges that the research community should address:
Perception of properties of novel objects: Develop advanced methods for accurately estimating key physical and semantic properties of new textiles. This challenge requires integrating sophisticated sensing technologies and exploratory procedures to identify these properties effectively. Achieving this would enhance robotic manipulation, enabling robots to adaptively handle a diverse range of textiles.
Multi-modal, Multi-task, Multi-embodiment Generalist Agent: Develop an agent capable of addressing multiple tasks in unknown environments, especially those involving the manipulation of deformable objects and assistive tasks with humans in the loop. This requires the agent to autonomously adapt to complex situations and acquire skills from minimal human demonstrations, leveraging multi-modal sensory inputs and versatile embodiment strategies. Achieving this would significantly advance robotic flexibility and effectiveness in diverse real-world applications.
Datasets and Benchmarking: Address the challenges in creating comprehensive datasets for high-dimensional state spaces and degrees of freedom in robotic manipulation. Develop a well-annotated dataset that captures 2D or 3D states across diverse, dynamic interactions labeled with relevant features and properties. Additionally, design standardized experimental benchmarking procedures for textile objects, considering stock availability and object variations, to facilitate consistent progress across research institutions. Achieving this will significantly enhance the community’s ability to develop and compare advanced robotic systems.